# MegaSaM: Accurate, Fast, and Robust Structure and Motion from Casual Dynamic Videos

### v1

@@ 4.1 Benchmarks and metrics
MPI Sintel. The MPI Sintel [4] dataset includes animated video sequences consisting of complex object motions and camera paths. Following CasualSAM [79], we evaluate all methods on the 18 sequences from the dataset, each of which consists of 20-50 images.
DyCheck. The DyCheck dataset [12] was initially designed for evaluating the task of novel view synthesis, and includes real-world videos of dynamic scenes captured from hand-held cameras. Each video includes 180-500 frames. We use the refined camera parameters and sensor depths provided by Shape of Motion [65] as ground truth.
@@
CasualSAM [79] 0.031 0.005 0.31 0.035 0.005 0.30 1.1m
LEAP-VO [6] 0.016 0.004 0.04 - - - 0.6s
ACE-Zero [3] 0.091 0.008 0.08 0.091 0.008 0.09 4.0s
Particle-SfM [80] 0.051 0.007 0.10 0.054 0.007 0.14 49s
RoDynRF [35] 0.116 0.021 0.34 0.112 0.031 0.39 7.6m
MonST3R [77] - - - 0.073 0.014 0.18 1.7s
@@
Sintel [4] Dycheck [12]
@@
DA-v2 [73] 0.37 0.55 58.6 0.20 0.27 84.7
DepthCrafter [20] 0.27 0.50 68.2 0.22 0.29 83.7
CasualSAM [79] 0.31 0.49 64.2 0.21 0.30 78.4
MonST3R [77] 0.31 0.43 62.5 0.26 0.35 66.5
@@
In-the-wild. We further evaluate on in-the-wild dynamic videos. Specifically, we include comparisons on 12 in-the-wild videos used by DynIBaR [33]. These videos feature long time duration (100-600 frames), uncontrolled camera paths, and complex scene motions. We construct ground truth movement masks via instance segmentation [17], where the instance IDs are manually specified, and use them to mask out moving objects before running COLMAP [52] to obtain reliable camera parameters.
@@
Sintel [4] Dycheck [12] In-the-Wild [33]
@@
Metrics. We use standard error metrics to evaluate camera pose estimation: Absolute Translation Error (ATE), Relative Translation Error (RTE), and Relative Rotation Error (RRE). Following CasualSaM, we normalize ground truth camera trajectories to be unit length, since camera tracks in different videos can vary significantly, and videos with longer trajectories can have higher influence on the computed metrics. For all methods, we align estimated camera paths to the ground truth trajectory by computing a global $\mathrm{Sim}(3)$ transformation via Umeyama alignment [63]. We report the average running time by dividing total running time of each method by the number of input frames. Additionally, we compare the quality of estimated video depths to recent baselines, adopting standard depth metrics: Absolute Relative Error, log RMSE, and Delta accuracy. We follow the standard evaluation protocol by excluding points that are further than 100 meters. For all methods, we align predicted video depths with ground truth through a global scale and shift estimate.
@@ 4.2 Quantitative Comparisons
Numerical results for camera pose estimation on the three benchmarks are reported in Tables 1, 2, and 3. Our approach demonstrates significant improvement and achieves the best camera tracking accuracy on all error metrics in both the calibrated and uncalibrated setting, while being competitive in terms of running time. Notably, our approach outperforms the concurrent work MonST3R [77] in terms of both robustness and accuracy even though MonST3R adopts a more recent global 3D point cloud representation for dynamic scenes. In addition, we report results for depth prediction on the Sintel and Dycheck in Table 4. Our depth estimates again outperform other baselines significantly in all metrics.
Figure 6: Visual comparisons of video depths. We compare video depth estimates from our approach and from CasualSAM [79] and MonST3R [77] by visualizing their depth maps (odd columns) and corresponding $x$-$t$ slices (even columns).
@@
Droid-SLAM [60] 0.030 0.022 0.50 - -
@@ 4.3 Ablation study
We perform an ablation study to validate major design choices for our camera tracking and depth estimation modules. Specifically, we evaluate camera tracking results with different configurations: 1) vanilla Droid-SLAM [60], 2) without mono-depth initialization (w/o mono-init.), 3) without object movement map prediction (w/o $\hat{m}_{i}$), 4) directly train the model on dynamic videos without using proposed two-stage training scheme (w/o 2-stage train.), 5) always turn on mono-depth regularization during global BA (w/o u-BA). We also ablate two major design decisions for video depth estimation: 1) jointly refine camera poses and depth estimation (w/o ft-pose), and 2) use original mono-depth prior loss from CasualSAM instead of our proposed one (w/o new $\mathcal{C}_{prior}$). As shown in Table 5, our full system outperforms all other alternative configurations.
Figure 7: We visualize two examples from DAVIS with rotation-dominant camera motion and narrow FoV. Compared to CasualSAM [79] and MonST3R [77], our system produces more accurate camera and geometry estimates of underlying dynamic scenes.
@@ 4.4 Qualitative Comparisons
Fig. 5 shows qualitative comparisons of estimated camera trajectory from our method and other baselines in 2D on three benchmarks, and our camera estimates are the closest to the ground truth. Furthermore, we visualize and compare the estimated video depths from our approach and two recent optimization based techniques, CasualSAM [79] and MonST3R [77], in Fig 6. In particular, we visualize both the depth map of a reference frame and corresponding $x$-$t$ depth slices over the entire video. Our approach once again produces more accurate, detailed and temporally consistent video depths.
@@
Last, we compare reconstruction and camera tracking quality among different approaches on challenging examples from DAVIS [43] by visualizing their estimated cameras and unprojected depth maps. As shown in Fig. 7, CasualSaM tends to produce distorted 3D point clouds, whereas MonST3R incorrectly treats rotational camera movement as translational one. In contrast, our approach produces more accurate camera along with more consistent geometry for such challenging inputs.
5 Discussion and conclusion
Limitations.
Despite excellent performance on a variety of in-the-wild videos, we observe that our approach can fail in extremely challenging scenarios, similar to findings from prior work [79]. For instance, camera tracking fails if moving objects dominate the entire image or if there is nothing for the system to track reliably. Please see supplemental material for visualization of failure cases. Furthermore, our system cannot handle videos with varying focal lengths or strong radial distortion within the video. Incorporating better priors from current vision foundation models into the pipelines is a promising future direction to explore.
Conclusion.
We presented a pipeline that produces accurate camera parameters and consistent depths from casual monocular videos of dynamic scenes. Our approach efficiently scales to in-the-wild footage of varying time duration, with unconstrained camera paths and complex scene dynamics. We have shown that, with careful extension, prior deep visual SLAM and SfM frameworks can be extended to achieve strong generalization to a broad range of videos and outperform recent state-of-the-art methods significantly.
Appendix A Implementation Details
Figure 9: System overview. Left: we estimate camera poses, focal length and low-resolution disparity maps from the input monocular video through differentiable Bundle Adjustment (BA): the network iteratively updates these state variables by learning to predict low-resolution flow $\hat{\mathbf{u}}_{ij}$, confidence, and movement probability maps $\mathbf{w}_{ij}$ and minimize weighted reprojection error between predicted flow ${\hat{\mathbf{u}}}_{ij}$ and flow induced by ego-motion ${\mathbf{u}}_{ij}$. We also initialize estimated disparity with mono-depth predicted from off-the-shelf models [72, 44]. Right: we fix estimated camera parameters and perform first-order global optimization over video depth and corresponding uncertainty parameters by minimizing flow and depth losses through pairwise 2D optical flows.
A.1 System Overview
Figure 9 shows an overview of our MegaSaM system. We separate the problem of camera and scene structure estimation into two stages, in the spirit of a conventional SfM pipeline [52, 53]. In particular, we first estimate camera poses $\hat{\mathbf{G}}$, focal length $\hat{f}$ and low-resolution disparity $\hat{\mathbf{d}}$ from the input monocular video through differentiable Bundle Adjustment (BA), where we initialize $\hat{\mathbf{d}}$ with monocular depth maps predicted from off-the-shelf models [72, 44]. In the second consistent video depth estimation phase, we fix estimated camera parameters and perform first-order optimization over video depth and uncertainty maps by enforcing flow and depth losses induced by pairwise 2D optical flows.
A.2 Framework and Architecture
We follow DROID-SLAM [59] for feature extraction, correlation feature construction, and perform iterative BA updates through flow, confidence, motion probability predictions. Each input to the model is a pair of video frames $(I_{i},I_{j})$.
Feature extraction. We use context and feature encoders to encode each input video frame into two different low-resolution feature maps at $\frac{1}{8}$ resolution of the input image, as shown in Figure 11.
Correlation feature construction. The correlation layer constructs a 4D correlation volume from the features encoded from an image pair, and each entry of the volume contains inner product of one pairs of feature vectors from the image pair.
Iterative updates. During each iterative BA step $k$, we update camera parameters and low-resolution disparity through flow, confidence and motion probability prediction. In particular, we first pretrain $F$ on synthetic video data (ego-motion pretraining in the main paper) to learn to predict flows and corresponding flow confidence, as shown by the gray blocks in Figure 10. In the second dynamic finetuning phase, we freeze the parameters of $F$ and finetune the motion module $F_{m}$ to predict extra object motion probability maps conditioned on the features from the ConvRGU, as shown in the blue blocks in Figure 10. Within the motion module, we first perform 2D spatial average pooling to provide the model with global spatial information; we then perform average pooling along the time axis to fuse information from $I_{i}$ and all its neighboring keyframes $I_{j}$ (where $j\in\mathcal{N}(i)$).
Figure 10: Architecture of flow, confidence and movement map predictor. The gray blocks belong to the network $F$ for flow and confidence prediction, and the blue blocks belong to the network $F_{m}$ for object movement map prediction. In the first stage, we perform ego-motion pretraining for $F$. In the second stage, we perform dynamic fine-tuning for $F_{m}$ while fixing the parameters of $F$.
Figure 11: Architecture of the feature and context encoders. Both encoders extract low-resolution features from input video frames at $\frac{1}{8}$ of the original resolution.
A.3 Consistent Video Depth Optimization
Recall, from Section 3.3 of our main paper, that we follow CasualSAM [79] to estimate consistent video depth by performing an additional first-order optimization on video disparity $\hat{D}_{i}$ along with per-frame aleatoric uncertainty maps $\hat{M}_{i}$. Instead of jointly optimizing camera parameters and scene structure as in CasualSAM, however, we fix camera parameters as done in conventional SfM pipelines like COLMAP [52, 53].
Our objective consists of three main cost functions:
We treat object motion in the video as the heteroscedastic aleatoric uncertainty of the flow reprojection and depth consistency error [23], and assume the underlying noise is Laplacian [74]. Specifically, for each selected pair $(I_{i},I_{j})$, flow reprojection loss $\mathcal{C}_{\text{flow}}$ compares $l_{1}$ loss weighted by the uncertainty $\hat{M}_{i}$ between flows $\text{flow}_{i\rightarrow j}$ from an off-the-shelf flow estimator [59] and the correspondences ${\mathbf{u}}_{ij}$ induced by our estimated camera motion and disparity through a multi-view constraint:
$\mathcal{C}^{i\rightarrow j}_{\text{flow}}$ $=\hat{M}_{i}||{\mathbf{u}}_{ij}-\mathbf{p}_{i},\text{flow}_{i\rightarrow j}(\mathbf{p}_{i})||_{1}+\log\left(\frac{1}{\hat{M}_{i}}\right),$ (13)
${{\mathbf{u}}_{ij}}$ $=\pi\left(\hat{\mathbf{G}}_{ij}\circ\pi^{-1}(\mathbf{p}_{i},\hat{D}_{i},K^{-1}),K\right)$ (14)
$\mathcal{C}_{\text{temp}}$ is an uncertainty weighted temporal depth loss that encourages pixel disparity to be temporally consistent according to estimated 2D optical flow:
$\mathcal{C}^{i\rightarrow j}_{\text{temp}}$ $=\hat{M}_{i}\delta\left(\mathbf{P}^{i\rightarrow j}_{z},\hat{D}_{j}(\mathbf{p}+\text{flow}_{i\rightarrow j}(\mathbf{p}))\right)+\log\left(\frac{1}{\hat{M}_{i}}\right)$
$\delta(a,b)$ $=||\max(\frac{a}{b},\frac{b}{a})||_{1}$
$\mathbf{P}^{i\rightarrow j}_{z}$ $=\left(D_{i}(\mathbf{p})\mathbf{R}_{i\rightarrow j}\mathbf{K}^{-1}\mathbf{p}+\mathbf{t}_{i\rightarrow j}\right)_{[z]}$ (15)
$\mathbf{R}_{i\rightarrow j}$ and $\mathbf{t}_{i\rightarrow j}$ are relative camera rotation and translation between $I_{i}$ and $I_{j}$; [z] is an operator that retrieve the third component of the 3D point vector (i.e. $z$ value).
$\mathcal{C}_{\text{prior}}$ is a depth prior loss that stops the final estimated video disparity from drifting too much from the initial estimate from the mono-depth network, and it consists of three losses:
$\mathcal{C}_{\text{prior}}=\mathcal{C}_{\text{si}}+w_{\text{grad}}\mathcal{C}_{\text{grad}}+w_{\text{normal}}\mathcal{C}_{\text{normal}}$ (16)
The scale-invariant depth loss $\mathcal{C}_{\text{si}}$ computes the mean square error of the difference among all pairs between optimized log-disparity $\log\hat{D}_{i}$ and initial log-disparity from the metric-aligned mono-depth prediction $\log D_{i}^{\text{align}}$.
$\mathcal{C}_{\text{si}}$ $=\frac{1}{n}\sum_{(\mathbf{p})}(R(\mathbf{p}))^{2}-\frac{1}{n^{2}}\left(\sum_{(\mathbf{p})}R(\mathbf{p})\right)^{2}$
$R_{i}$ $=\log(\hat{D}_{i})-\log(D_{i}^{\text{align}}).$ (17)
$\mathcal{C}_{\text{grad}}$ is a multi-scale scale-invariant gradient matching term [30], which computes $l_{1}$ difference between estimated log disparity gradients and initial log-disparity gradients
$\mathcal{C}_{\text{grad}}$ $=\frac{1}{n}\sum_{s}w_{\nabla}^{s}(\mathbf{p})\sum_{\mathbf{p}}\left(|\nabla_{x}R^{s}(\mathbf{p})|+|\nabla_{y}R^{s}(\mathbf{p})|\right)$
$w_{\nabla}^{s}(\mathbf{p})$ $=1-\exp\left(-\beta_{\nabla}(\nabla_{x}R^{s}(\mathbf{p})+\nabla_{y}R^{s}(\mathbf{p}))\right)$ (18)
where $R^{s}(\mathbf{p})$ is log-depth difference map at pixel position $\mathbf{p}$ and scale $s$. In other words, we only apply multi-scale gradient matching loss to pixels where the current estimated disparity deviates significantly from the original mono-depth.
$\mathcal{C}_{\text{normal}}$ is a surface normal loss that encourages that normal $\hat{\mathbf{N}}(\mathbf{p})$ derived from estimated disparity to be close to the surface normal $\mathbf{N}^{\text{align}}$ derived from the initial metric-aligned monocular disparity:
$\mathcal{C}_{\text{normal}}=\sum_{\mathbf{p}}1-\hat{\mathbf{N}}(\mathbf{p})\cdot\mathbf{N}^{\text{align}}(\mathbf{p})$ (19)
We set $w_{\text{grad}}=1,w_{\text{normal}}=4,\beta_{\nabla}=5$ throughout our experiments. We simply choose image pairs $(I_{i},I_{j})$ from a set of fixed intervals following prior work [79]: $j\in(i+1,i+2,i+4,i+8,i+15)$. During optimization, we initialize the disparity variables from the metric-aligned monocular depth by combining estimates from off-the-shelf modules as described in the main paper [72, 44], and we initialize the uncertainty map with object motion probability maps predicted from our camera tracking module. The optimization first conducts a “warm-up” phase for 100 steps by fixing the video disparity variables and optimizing the per-frame uncertainty map, per-frame scale, shift variables using the aforementioned losses. The disparity maps and uncertainty maps are then optimized together under the aforementioned losses for another 400 steps.
A.4 Additional Details
Training Losses. We supervise our network using a combination of pose loss and flow loss. The flow loss is applied to pairs of adjacent frames. We compute the optical flow induced by the predicted depth and poses and the flow induced by the ground truth depth and poses. The loss is taken to be the average l2 distance between the two flow fields.
Given a set of ground truth poses $\{\mathbf{T}_{i}\}_{i=1}^{N}$ and predicted poses $\{\mathbf{G}_{i}\}_{i=1}^{N}$, the pose loss is taken to be the distance between the ground truth and predicted poses, $\mathcal{L}_{pose}=\sum_{i}||\text{Log}_{SE(3)}(\mathbf{T}_{i}^{-1}\cdot\mathbf{G}_{i})||_{2}$. We apply the losses to the output of every BA iteration with exponentially increasing weight using $\gamma=0.9^{k}$, where $k$ indicates the $k^{th}$ BA iterations.
Training and Inference Details In our two-stage training scheme, we first pretrain our model on synthetic data of static scenes, which include 163 scenes from TartanAir [69] and 5K videos from static Kubric [15]. In the second stage, we finetune motion module $F_{m}$ on 11K dynamic videos from Kubric [15]. Each training example consists of a 7-frame video sequence. We first precompute a distance matrix between each pair of video frame based on the average ego-motion induced flow magnitude. We then dynamically generate a training sequence according to the constructed distance matrix, we randomly sample each frame such that average flow between them is between 0.5px and 64px.
Within the camera tracking module, we normalize video disparity $\hat{\mathbf{d}}$ such that its $98$ percentile is $2$; we also normalize focal length by dividing it by the input image resolution within every the bundle adjustment stage.
Appendix B Limitations
Despite excellent performance on a variety of in-the-wild videos, we observe that our approach can fail in extremely challenging scenarios, similar to findings from prior work [79]. For instance, camera tracking fails if moving objects dominate the entire image or if there is nothing for the system to track reliably, as shown in the first row of Fig. 8. Furthermore, our approach also struggles on dynamic videos where camera motion and object motion are colinear, as shown in the second row of Fig. 8.

### v2

@@ 4.1 Benchmarks and metrics
MPI Sintel. The MPI Sintel [Butler:ECCV:2012] dataset includes animated video sequences consisting of complex object motions and camera paths. Following CasualSAM [zhang2022structure], we evaluate all methods on the 18 sequences from the dataset, each of which consists of 20-50 images.
DyCheck. The DyCheck dataset [gao2022monocular] was initially designed for evaluating the task of novel view synthesis, and includes real-world videos of dynamic scenes captured from hand-held cameras. Each video includes 180-500 frames. We use the refined camera parameters and sensor depths provided by Shape of Motion [wang2024shape] as ground truth.
@@
CasualSAM [zhang2022structure] 0.031 0.005 0.31 0.035 0.005 0.30 1.1m
LEAP-VO [chen2024leap] 0.016 0.004 0.04 - - - 0.6s
ACE-Zero [brachmann2024scene] 0.091 0.008 0.08 0.091 0.008 0.09 4.0s
Particle-SfM [zhao2022particlesfm] 0.051 0.007 0.10 0.054 0.007 0.14 49s
RoDynRF [liu2023robust] 0.116 0.021 0.34 0.112 0.031 0.39 7.6m
MonST3R [zhang2024monst3r] - - - 0.073 0.014 0.18 1.7s
@@
Sintel [Butler:ECCV:2012] Dycheck [gao2022monocular]
@@
DA-v2 [yang2024depth] 0.37 0.55 58.6 0.20 0.27 84.7
DepthCrafter [hu2024depthcrafter] 0.27 0.50 68.2 0.22 0.29 83.7
CasualSAM [zhang2022structure] 0.31 0.49 64.2 0.21 0.30 78.4
MonST3R [zhang2024monst3r] 0.31 0.43 62.5 0.26 0.35 66.5
@@
In-the-wild. We further evaluate on in-the-wild dynamic videos. Specifically, we include comparisons on 12 in-the-wild videos used by DynIBaR [li2023dynibar]. These videos feature long time duration (100-600 frames), uncontrolled camera paths, and complex scene motions. We construct ground truth movement masks via instance segmentation [he2017mask], where the instance IDs are manually specified, and use them to mask out moving objects before running COLMAP [schonberger2016structure] to obtain reliable camera parameters.
@@
Sintel [Butler:ECCV:2012] Dycheck [gao2022monocular] In-the-Wild [li2023dynibar]
@@
Metrics. We use standard error metrics to evaluate camera pose estimation: Absolute Translation Error (ATE), Relative Translation Error (RTE), and Relative Rotation Error (RRE). Following CasualSaM, we normalize ground truth camera trajectories to be unit length, since camera tracks in different videos can vary significantly, and videos with longer trajectories can have higher influence on the computed metrics. For all methods, we align estimated camera paths to the ground truth trajectory by computing a global $\mathrm{Sim}(3)$ transformation via Umeyama alignment [umeyama1991least]. We report the average running time by dividing total running time of each method by the number of input frames. Additionally, we compare the quality of estimated video depths to recent baselines, adopting standard depth metrics: Absolute Relative Error, log RMSE, and Delta accuracy. We follow the standard evaluation protocol by excluding points that are further than 100 meters. For all methods, we align predicted video depths with ground truth through a global scale and shift estimate.
@@ 4.2 Quantitative Comparisons
Numerical results for camera pose estimation on the three benchmarks are reported in Tables 1, 2, and 3. Our approach demonstrates significant improvement and achieves the best camera tracking accuracy on all error metrics in both the calibrated and uncalibrated setting, while being competitive in terms of running time. Notably, our approach outperforms the concurrent work MonST3R [zhang2024monst3r] in terms of both robustness and accuracy even though MonST3R adopts a more recent global 3D point cloud representation for dynamic scenes. In addition, we report results for depth prediction on the Sintel and Dycheck in Table 4. Our depth estimates again outperform other baselines significantly in all metrics.
Figure 6: Visual comparisons of video depths. We compare video depth estimates from our approach and from CasualSAM [zhang2022structure] and MonST3R [zhang2024monst3r] by visualizing their depth maps (odd columns) and corresponding $x$-$t$ slices (even columns).
@@
Droid-SLAM [teed2021droid] 0.030 0.022 0.50 - -
@@ 4.3 Ablation study
We perform an ablation study to validate major design choices for our camera tracking and depth estimation modules. Specifically, we evaluate camera tracking results with different configurations: 1) vanilla Droid-SLAM [teed2021droid], 2) without mono-depth initialization (w/o mono-init.), 3) without object movement map prediction (w/o $\hat{m}_{i}$), 4) directly train the model on dynamic videos without using proposed two-stage training scheme (w/o 2-stage train.), 5) always turn on mono-depth regularization during global BA (w/o u-BA). We also ablate two major design decisions for video depth estimation: 1) jointly refine camera poses and depth estimation (w/o ft-pose), and 2) use original mono-depth prior loss from CasualSAM instead of our proposed one (w/o new $\mathcal{C}_{prior}$). As shown in Table 5, our full system outperforms all other alternative configurations.
Figure 7: We visualize two examples from DAVIS with rotation-dominant camera motion and narrow FoV. Compared to CasualSAM [zhang2022structure] and MonST3R [zhang2024monst3r], our system produces more accurate camera and geometry estimates of underlying dynamic scenes.
@@ 4.4 Qualitative Comparisons
Fig. 5 shows qualitative comparisons of estimated camera trajectory from our method and other baselines in 2D on three benchmarks, and our camera estimates are the closest to the ground truth. Furthermore, we visualize and compare the estimated video depths from our approach and two recent optimization based techniques, CasualSAM [zhang2022structure] and MonST3R [zhang2024monst3r], in Fig 6. In particular, we visualize both the depth map of a reference frame and corresponding $x$-$t$ depth slices over the entire video. Our approach once again produces more accurate, detailed and temporally consistent video depths.
@@
Last, we compare reconstruction and camera tracking quality among different approaches on challenging examples from DAVIS [Perazzi2016] by visualizing their estimated cameras and unprojected depth maps. As shown in Fig. 7, CasualSaM tends to produce distorted 3D point clouds, whereas MonST3R incorrectly treats rotational camera movement as translational one. In contrast, our approach produces more accurate camera along with more consistent geometry for such challenging inputs.